if an investment of $14,000 grows to $15,300 in 21 years, what is the interest rate assuming quaterly compounding.
your final answer must be in percent, accurate to 4 decimal places.



B. what is the interest rate if we use a simple interest?
your final answer must be in percent, accurate to 4 decimal places

Answer by stanbon(75887)

if an investment of $14,000 grows to $15,300 in 21 years, what is the interest rate assuming quaterly compounding.
your final answer must be in percent, accurate to 4 decimal places.
A=P(1+r/n)^(nt)
15,300=14,000(1+r/4)(21*4)
1.092857143.. = (1+r/4)^(44)
Take the 44th root of both sides to get:
1.002020117=1+(r/4)
0.002020117=r/4
r=0.00808046881
rate = 0.8080%
B. what is the interest rate if we use a simple interest?
your final answer must be in percent, accurate to 4 decimal places
A=P(1+rt)^n
15,300=14000(1+r*4)^21
1.092857143=(1+r/4)^21
Take the 21st root of both sides to get:
1.004237309 = 1+r/4
0.004237309=r/4
r=0.169492369
rate = 1.695%
Cheers,
Stan H.